# EvoGPT: Leveraging LLM-Driven Seed Diversity to Improve Search-Based Test Suite Generation

### v2

@@ I Introduction
Software testing is a fundamental practice in software engineering, critical for ensuring the reliability, correctness, and maintainability of software systems [beck2003tdd]. Among its various forms, unit testing serves as the first line of defense against bugs by verifying the behavior of individual program units in isolation. However, despite its importance, the process of manually designing high-quality unit tests remains labor-intensive and error-prone.
@@
Search-Based Software Testing (SBST) is an extensively studied paradigm for automated test generation [harman2007searchbased, almulla2022learning, lukasczyk2022pynguin]. Techniques such as Evolutionary Algorithms (EAs), implemented in tools like EvoSuite [fraser2012evosuite], and DynaMosa[panichella2017automated] systematically optimize test suites toward structural coverage objectives (e.g., branch and line coverage). While effective, SBST techniques are known to suffer from limitations such as coverage plateaus - where evolutionary search stagnates despite additional generations, and difficulty generating semantically diverse tests for complex behaviors.
@@
In recent years, Large Language Models (LLMs) have emerged as promising tools for automating unit test generation [chen2021codex]. Approaches such as ChatUniTest [chen2024chatunitest], TestART [2024arXiv240803095G], and A3Test [xie2024a3test] demonstrate the ability of LLMs to produce syntactically correct, semantically coherent, and human-readable unit tests. These approaches leverage their natural language understanding and code generation capabilities to approximate human-level testing strategies. Nevertheless, LLM-based methods often suffer from limited exploration of complex input spaces, instability caused by hallucinations, and an inability to iteratively improve upon initial generations without external feedback [madaan2023self].
@@
Motivated by the complementary strengths and weaknesses of SBST and LLM-based approaches, we propose EvoGPT, a hybrid system that integrates diverse LLM-based test generation with evolutionary search-based test suite optimization. EvoGPT is illustrated in Figure 1. It generates a diverse initial population of test suites by querying LLMs under multiple system prompts and temperature settings, encouraging exploration across different execution paths and behavioral patterns. A generation-repair loop and a coverage enhancement step inspired by TestART [2024arXiv240803095G] further refines this initial population. Subsequently, EvoGPT uses an Evolutionary Algorithm (EA) that applies selection, crossover, and mutation operators to optimize a fitness function that combines structural coverage and mutation score.
@@
A key insight underlying EvoGPT is that the effectiveness of LLM-SBST hybridization critically depends on injecting semantically diverse test suites at the right stages of the evolutionary optimization process, rather than relying on a single-shot LLM initialization (i.e., generating a single LLM-based test population once, without further LLM involvement during search) or uniform intervention throughout the search. Accordingly, EvoGPT incorporates a plateau escape mechanism inspired by CodaMOSA [lemieux2023codamosa]: when the evolutionary optimization process stagnates i.e., no significant fitness improvement is observed for a configurable number of generations, EvoGPT re-invokes multiple prompt and temperature LLM configurations to generate targeted test methods aimed at uncovered branches and mutations. These tests are injected into the current best test suite, enabling the search to escape local optima and continue improving.
@@
EvoGPT is not the first test generation system to integrate search-based and LLM-based test generation methods [xiao2024optimizing, dakhama2025enhancing, lemieux2023codamosa, yang2025llm]. Most prior work focused on the question of when LLMs should be used in SBST, showing they are effective during initialization and after the search reached a plateau [xiao2024optimizing, lemieux2023codamosa]. In this work, we focus on a complementary question: how to effectively generate and exploit semantically diverse LLM-generated candidates through prompt and temperature diversity, both at initialization and during plateau escaping, under equal evolutionary budgets.
@@
We evaluate EvoGPT on the Defects4J benchmark [just2014defects4j] and compare it against state-of-the-art SBST-only and LLM-only baselines, namely EvoSuite [fraser2012evosuite] and TestART [2024arXiv240803095G]. Our results show that EvoGPT significantly outperforms both baselines in line coverage, branch coverage, and mutation score of successfully generated tests. An ablation study further demonstrates the contribution of prompt diversity, evolutionary optimization, and plateau escape to EvoGPT’s overall effectiveness.
@@
  • Empirical evaluation. Experimental results using the Defects4J benchmark [just2014defects4j] show the benefits of using our EvoGPT system, demonstrating substantial improvements in code coverage and mutation testing over strong baselines, namely EvoSuite [fraser2012evosuite] and TestART [2024arXiv240803095G], with gains driven by explicitly enforcing diversity in the generated test suites.
@@ II Background and Problem Definition
A unit test is a self-contained function designed to validate the behavior of a specific unit of source code, most commonly an individual method. Executing a test typically involves invoking the target method under specific inputs and comparing the observed outcome against expected outputs through assertions. The result of a test is often binary: a test either passes or fails, [^1] In practice, tests may also exhibit flakiness, where outcomes vary across executions due to non-determinism, environmental dependencies, or timing effects. In this work, we focus on deterministic unit tests and consider only consistently passing tests when evaluating test quality. depending on whether all assertions hold. The objective of an automated unit test generation algorithm is to synthesize a suite of correct, high-quality, unit tests. Following prior work [2024arXiv240803095G], we designate the public methods of each public class under test as focal methods, which constitute the primary targets for test generation and evaluation. This focus enables the testing effort to concentrate on actionable and externally observable behavior. The following metrics are commonly used for evaluating the quality of a suite of unit tests [fraser2012evosuite, harman2007searchbased, jia2011survey]:
@@
  • Mutation Score of Correct Tests (MSCT): the percentage of artificially injected faults (mutants) in the focal methods that are detected, i.e., cause at least one test to fail. This is a commonly used proxy for fault detection capability [fraser2012evosuite, jia2011survey].
@@
Many techniques have been proposed to automatically generate test suites. Random test generation methods such as Randoop [pacheco2007randoop] sample inputs and assertions possibly applying heuristics for guidance. Assertions are inferred from observed runtime behavior, for example by recording returned values, object states, or thrown exceptions, and then checking for consistency in subsequent executions. These methods are computationally inexpensive, but often produce shallow tests with limited coverage and weak or missing assertions.
Search-Based Software Testing (SBST) methods, such as EvoSuite [fraser2012evosuite] apply evolutionary algorithms to optimize test quality metrics such as coverage and mutation score over multiple generations. These SBST methods begin by generating an initial population of random test suites, which are then iteratively refined using three core operators: selection, which favors test suites with higher fitness for reproduction; crossover, which combines parts of two parent test suites to create new ones; and mutation, which introduces small random changes to individual tests such as modifying numeric or string literals within the test, removing assertions, etc. This evolutionary process gradually improves coverage and fault detection ability over time. Notably, applying evolutionary algorithms often requires implementing domain-specific fitness functions, cross-over and mutation operators. While EvoSuite and similar approaches can achieve high structural coverage, they struggle with assertion quality and test readability [biagiola2025improving].
Other methods for test generation require either additional knowledge of the software being tested or the ability to execute and instrument the tests to trace their execution. For example, metamorphic test generation assume as input a set of properties that relate the values of the tested units’ inputs and outputs. These properties, known as metamorphic relations, guide the generation of follow-up tests whose expected behavior can be inferred from previous executions [chen2018metamorphic]. Test generation methods based on dynamic symbolic execution [chen2013state] execute a program while simultaneously tracking symbolic expressions over input variables, capturing the path conditions encountered during execution. These path constraints are then solved using constraint solvers (e.g., SMT solvers) to generate new inputs that drive execution toward previously unexplored branches.
@@
Recently, LLMs have been used to synthesize test cases based on natural language prompts or code structure. Early approaches relied on single shot prompts (e.g., “Write a unit test for this method”) [yuan2024chatgpt], but more sophisticated techniques have since emerged. These include prompt engineering to guide assertion behavior, few-shot prompting to improve test structure [chen2024chatunitest], and iterative refinement using feedback such as stack traces or compilation errors[2024arXiv240803095G].
For example, TestART [2024arXiv240803095G] combines LLM-based test generation with a feedback loop that repairs failing tests. If a generated test fails to compile or run, it captures the error and tries to fix it programmatically using a set of repair rules. If the error persists, it re-prompts the model with the error trace and original input, refining the test in multiple rounds. Meta’s Automated Compliance Hardening (ACH) [harman2025mutation] is another example of an LLM-based test generation system. They focus on generating tests designed to find a specific type of bugs, using an LLM to generate mutations that simulate such bugs and then to generate tests that catch them.
LLM-generated tests, however, often contain superficial validations and tend to overfit syntactic patterns rather than exercising the underlying semantics of the code under test [yuan2024chatgpt, chen2024chatunitest, testpilot2023]. Empirical comparisons have shown that, while LLM-based approaches can achieve reasonable mutation scores, they can lag behind search-based and symbolic execution techniques in structural coverage and could be sensitive to the complexity of the code under test [abdullin2025test]. Moreover, the assertions generated by LLMs are frequently shallow, avoid exercising deep program logic, and fail to substantially improve test quality in the absence of external feedback.
@@ II-C LLMs and Evolutionary Algorithms
Recent works have begun to explore how LLMs and Evolutionary Algorithms (EAs) can complement one another for test generation and for other tasks. Broadly, LLMs can (i) act as evolutionary operators by proposing candidate solutions like seeding the initial population, (ii) provide semantic guidance that helps EAs escape local optima, or (iii) encode problem structure that would otherwise be hard to capture with hand-designed heuristics. Liu et al. [liu2024large] showed that an LLM can be prompted to perform selection, crossover, and mutation directly, demonstrating that LLMs can function as zero-shot evolutionary optimizers. Wang et al. [wang2025multi] further extended this idea to multi-agent settings where specialized LLM agents collectively drive an evolutionary search process.
Wu et al. [wu2024evolutionary] provide a taxonomy of these approaches, distinguishing between LLM-enhanced EAs (using LLMs to support or improve evolutionary search), EA-enhanced LLMs (using evolutionary strategies to tune or guide LLMs), and tightly integrated hybrid systems. EvoGPT fits into the category of LLM-enhanced EAs. Rather than replacing the evolutionary loop, we use LLMs to generate a semantically diverse initial population and to provide guided injections when the search plateaus. This combination retains the strengths of traditional SBST while exploiting the generative and semantic capabilities of modern LLMs.
@@ II-D LLMs and EAs for Automated Test Generation
The integration of LLMs and EAs specifically for automated test generation has recently gained momentum, motivated by the complementary strengths of semantic reasoning from LLMs and systematic exploration from EAs. Xiao et al. [xiao2024optimizing] investigate three places where LLM assistance can be used in an EA for test generation: (i) generating the initial population of test suites, (ii) injecting additional test cases during the evolutionary process to target uncovered code, and (iii) introducing new candidates when the search stagnates. Their experimental results show that LLM-based assistance yields the strongest benefits during the initialization phase and when applied after moderate stagnation, whereas overly frequent intervention provides diminishing returns.
Another influential system in this space is CodaMosa by Microsoft Research [lemieux2023codamosa]. CodaMosa is an EA-based software test generation system that uses an LLM to escape structural coverage plateaus. Rather than relying solely on mutation and crossover, CodaMosa periodically queries an LLM to synthesize new tests aimed towards areas of the tested program that are less covered. We implemented a similar plateau-escaping mechanism in our EvoGPT.
Yang et al. [yang2025llm] created pytLMtester, a test generation system that extends DynaMosa [panichella2017automated] for dynamically typed languages using LLMs. They run an EA to generate tests and use LLMs to infer object types, repair failed tests, and generate mutations. Hybrid methods for automated test generation have also been explored in other domains. SearchSYS [dakhama2025enhancing] uses LLMs to generate high-level test templates that guide a search-based fuzzer in system-level simulators, showing that LLM-derived structure can improve bug-finding effectiveness.
@@ III The EvoGPT System
The initial population generation process instantiates five distinct LLM agent configurations, each defined by a different prompting strategy (e.g., focusing on edge cases, deep object chains, or creative assertion styles) as well as different temperature values. Each agent is queried asynchronously five times, resulting in a total of 25 diverse and structurally unique test suites. During this generation process, a lightweight repair loop inspired by TestART [2024arXiv240803095G], is applied to each test suite: when a generated test fails due to compilation or runtime errors, the system captures the stack trace and attempts to fix it. This helps improve test validity and robustness early in the pipeline. Successful test suites go through an additional coverage enhancement step, which involves identifying uncovered branches in the code and calling an LLM to attempt to add tests that cover them. The resulting 25 test suites are passed to EvoGPT’s population optimization process, in which an EA iteratively evolves this population by applying selection, crossover, and mutation operations guided by a fitness function. Inspired by CodaMosa [lemieux2023codamosa], when EvoGPT detects stagnation in the EA, it injects additional tests to help escape local optima. Here as well, multiple LLM agents are used to generate these injected tests, to encourage diversity. Next, we describe the details of the EvoGPT main processes: initial population generation process and population optimization.
@@ III-A Initial Population Generation
Prior to starting the initial population generation process, EvoGPT performs a lightweight pre-processing on the source code to reduce token complexity and context window. Specifically, we remove excessive or lengthy inline comments, documentation blocks, and unreachable code segments that could overload the model’s context window and negatively affect generation quality [gao-etal-2024-insights]. Then, we generate 25 test suites for a given class by instantiating five LLM agents and having each agent generate five test suites.
Importantly, each LLM agent is configured with a different temperature setting and a corresponding system prompt. The temperature value is used to control the stochasticity of the language model output. Prior research has shown that varying temperature values leads to a broader range of generated outputs, exposing different behavioral patterns and increasing semantic variance across samples [agarwal2024temperature]. To complement this, each agent is given a custom system prompt, which directs it to approach the test suite generation task from a slightly different perspective (e.g., focusing on execution paths, edge values, or assertion robustness). Table I lists for each of the five types of LLM agents we used, their temperature value, a qualitative description of their prompt (column “Prompt objective”), and intended purpose (column “Purpose”). The exact system prompts used for each LLM agent are included in the provided code. This diversity promotes both semantic variability and robustness in the generated tests. We empirically verified that this multi-prompt configuration produces semantically diverse initial test suites. A detailed analysis of the diversity metrics and statistical tests is presented in Section IV.
@@
After generating the initial set of tests, each LLM agent performs a generation-repair loop where it iteratively refines its test case in response to runtime feedback such as compilation errors or failing assertions. Such a generation-repair process was shown to enhance test validity and robustness in prior LLM-based systems such as ChatUniTest [chen2024chatunitest], TestART [2024arXiv240803095G] and TestPilot [testpilot2023]. In our system, this loop leverages a simplified version of TestART’s repair strategy: when a generated test fails with a runtime error or stack trace, EvoGPT programmatically extracts the stack trace and invokes a minimal repair heuristic to remove or correct the failing statement. Our minimal repair heuristic directly fixes common runtime and compilation errors. For instance, when an import error is raised, the repair loop programmatically traverses through the source code’s repository in order to fetch the right import paths. If our minimal repair heuristic fails to fix the generated test, we call the LLM with a repair prompt along with the stack trace that define the runtime or compilation error. The exact repair prompt is also included in our code. Using a programmatic, targeted, template-based repairs before calling an LLM-based repair has proved to be more stable and cost-effective for improving test pass rates than only using an LLM-based repair [2024arXiv240803095G]
This repair loop is repeated at most four times, bounding the accumulation of error traces and code context passed to the LLM, thereby limiting prompt growth and reducing the risk of context window saturation [gao-etal-2024-insights]. If errors in the tests persist, the test methods responsible for these errors are automatically removed from the test suite.
@@
Once a valid test suite is synthesized, it is compiled and executed using JaCoCo,222www.jacoco.org/jacoco a Java Code Coverage tool, which provides detailed line and branch coverage information. The resulting coverage report is parsed and passed to the next module: the Coverage Analysis Agent. This agent uses the same temperature as the original generation agent to maintain behavioral consistency. It is then instructed to generate additional complementary tests, aimed specifically at covering the missed branches and lines reported in the coverage report. This strategy was found valuable for enhancing code coverage, as demonstrated in TestART [2024arXiv240803095G]. Finally, the tests produced by both the initial generation agent and its corresponding coverage enhancement process are merged into a single composite test suite. This forms one member of the initial population in our evolutionary algorithm.

### v3

@@ I Introduction
Software testing is a fundamental practice in software engineering, critical for ensuring the reliability, correctness, and maintainability of software systems [7]. Among its various forms, unit testing serves as the first line of defense against bugs by verifying the behavior of individual program units in isolation. However, despite its importance, the process of manually designing high-quality unit tests remains labor-intensive and error-prone.
@@
Search-Based Software Testing (SBST) is an extensively studied paradigm for automated test generation [20, 4, 25]. Techniques such as Evolutionary Algorithms (EAs), implemented in tools like EvoSuite [15], and DynaMosa[28] systematically optimize test suites toward structural coverage objectives (e.g., branch and line coverage). While effective, SBST techniques are known to suffer from limitations such as coverage plateaus - where evolutionary search stagnates despite additional generations, and difficulty generating semantically diverse tests for complex behaviors.
@@
In recent years, Large Language Models (LLMs) have emerged as promising tools for automating unit test generation [9]. Approaches such as ChatUniTest [12], TestART [18], and A3Test [3] demonstrate the ability of LLMs to produce syntactically correct, semantically coherent, and human-readable unit tests. These approaches leverage their natural language understanding and code generation capabilities to approximate human-level testing strategies. Nevertheless, LLM-based methods often suffer from limited exploration of complex input spaces, instability caused by hallucinations, and an inability to iteratively improve upon initial generations without external feedback [26].
@@
Motivated by the complementary strengths and weaknesses of SBST and LLM-based approaches, we propose EvoGPT, a hybrid system that integrates diverse LLM-based test generation with evolutionary search-based test suite optimization. EvoGPT is illustrated in Figure 1. It generates a diverse initial population of test suites by querying LLMs under multiple system prompts and temperature settings, encouraging exploration across different execution paths and behavioral patterns. A generation-repair loop and a coverage enhancement step inspired by TestART [18] further refines this initial population. Subsequently, EvoGPT uses an Evolutionary Algorithm (EA) that applies selection, crossover, and mutation operators to optimize a fitness function that combines structural coverage and mutation score.
@@
A key insight underlying EvoGPT is that the effectiveness of LLM-SBST hybridization critically depends on injecting semantically diverse test suites at the right stages of the evolutionary optimization process, rather than relying on a single-shot LLM initialization (i.e., generating a single LLM-based test population once, without further LLM involvement during search) or uniform intervention throughout the search. Accordingly, EvoGPT incorporates a plateau escape mechanism inspired by CodaMOSA [23]: when the evolutionary optimization process stagnates i.e., no significant fitness improvement is observed for a configurable number of generations, EvoGPT re-invokes multiple prompt and temperature LLM configurations to generate targeted test methods aimed at uncovered branches and mutations. These tests are injected into the current best test suite, enabling the search to escape local optima and continue improving.
@@
EvoGPT is not the first test generation system to integrate search-based and LLM-based test generation methods [34, 14, 23, 35]. Most prior work focused on the question of when LLMs should be used in SBST, showing they are effective during initialization and after the search reached a plateau [34, 23]. In this work, we focus on a complementary question: how to effectively generate and exploit semantically diverse LLM-generated candidates through prompt and temperature diversity, both at initialization and during plateau escaping, under equal evolutionary budgets.
@@
We evaluate EvoGPT on the Defects4J benchmark [22] and compare it against state-of-the-art SBST-only and LLM-only baselines, namely EvoSuite [15] and TestART [18]. Our results show that EvoGPT significantly outperforms both baselines in line coverage, branch coverage, and mutation score of successfully generated tests. An ablation study further demonstrates the contribution of prompt diversity, evolutionary optimization, and plateau escape to EvoGPT’s overall effectiveness.
@@
  • Empirical evaluation. Experimental results using the Defects4J benchmark [22] show the benefits of using our EvoGPT system, demonstrating substantial improvements in code coverage and mutation testing over strong baselines, namely EvoSuite [15] and TestART [18], with gains driven by explicitly enforcing diversity in the generated test suites.
@@ II Background and Problem Definition
A unit test is a self-contained function designed to validate the behavior of a specific unit of source code, most commonly an individual method. Executing a test typically involves invoking the target method under specific inputs and comparing the observed outcome against expected outputs through assertions. The result of a test is often binary: a test either passes or fails, [^1] In practice, tests may also exhibit flakiness, where outcomes vary across executions due to non-determinism, environmental dependencies, or timing effects. In this work, we focus on deterministic unit tests and consider only consistently passing tests when evaluating test quality. depending on whether all assertions hold. The objective of an automated unit test generation algorithm is to synthesize a suite of correct, high-quality, unit tests. Following prior work [18], we designate the public methods of each public class under test as focal methods, which constitute the primary targets for test generation and evaluation. This focus enables the testing effort to concentrate on actionable and externally observable behavior. The following metrics are commonly used for evaluating the quality of a suite of unit tests [15, 20, 21]:
@@
  • Mutation Score of Correct Tests (MSCT): the percentage of artificially injected faults (mutants) in the focal methods that are detected, i.e., cause at least one test to fail. This is a commonly used proxy for fault detection capability [15, 21].
@@
Many techniques have been proposed to automatically generate test suites. Random test generation methods such as Randoop [27] sample inputs and assertions possibly applying heuristics for guidance. Assertions are inferred from observed runtime behavior, for example by recording returned values, object states, or thrown exceptions, and then checking for consistency in subsequent executions. These methods are computationally inexpensive, but often produce shallow tests with limited coverage and weak or missing assertions.
Search-Based Software Testing (SBST) methods, such as EvoSuite [15] apply evolutionary algorithms to optimize test quality metrics such as coverage and mutation score over multiple generations. These SBST methods begin by generating an initial population of random test suites, which are then iteratively refined using three core operators: selection, which favors test suites with higher fitness for reproduction; crossover, which combines parts of two parent test suites to create new ones; and mutation, which introduces small random changes to individual tests such as modifying numeric or string literals within the test, removing assertions, etc. This evolutionary process gradually improves coverage and fault detection ability over time. Notably, applying evolutionary algorithms often requires implementing domain-specific fitness functions, cross-over and mutation operators. While EvoSuite and similar approaches can achieve high structural coverage, they struggle with assertion quality and test readability [8].
Other methods for test generation require either additional knowledge of the software being tested or the ability to execute and instrument the tests to trace their execution. For example, metamorphic test generation assume as input a set of properties that relate the values of the tested units’ inputs and outputs. These properties, known as metamorphic relations, guide the generation of follow-up tests whose expected behavior can be inferred from previous executions [11]. Test generation methods based on dynamic symbolic execution [10] execute a program while simultaneously tracking symbolic expressions over input variables, capturing the path conditions encountered during execution. These path constraints are then solved using constraint solvers (e.g., SMT solvers) to generate new inputs that drive execution toward previously unexplored branches.
@@
Recently, LLMs have been used to synthesize test cases based on natural language prompts or code structure. Early approaches relied on single shot prompts (e.g., “Write a unit test for this method”) [36], but more sophisticated techniques have since emerged. These include prompt engineering to guide assertion behavior, few-shot prompting to improve test structure [12], and iterative refinement using feedback such as stack traces or compilation errors[18].
For example, TestART [18] combines LLM-based test generation with a feedback loop that repairs failing tests. If a generated test fails to compile or run, it captures the error and tries to fix it programmatically using a set of repair rules. If the error persists, it re-prompts the model with the error trace and original input, refining the test in multiple rounds. Meta’s Automated Compliance Hardening (ACH) [19] is another example of an LLM-based test generation system. They focus on generating tests designed to find a specific type of bugs, using an LLM to generate mutations that simulate such bugs and then to generate tests that catch them.
LLM-generated tests, however, often contain superficial validations and tend to overfit syntactic patterns rather than exercising the underlying semantics of the code under test [36, 12, 29]. Empirical comparisons have shown that, while LLM-based approaches can achieve reasonable mutation scores, they can lag behind search-based and symbolic execution techniques in structural coverage and could be sensitive to the complexity of the code under test [1]. Moreover, the assertions generated by LLMs are frequently shallow, avoid exercising deep program logic, and fail to substantially improve test quality in the absence of external feedback.
@@ II-C LLMs and Evolutionary Algorithms
Recent works have begun to explore how LLMs and Evolutionary Algorithms (EAs) can complement one another for test generation and for other tasks. Broadly, LLMs can (i) act as evolutionary operators by proposing candidate solutions like seeding the initial population, (ii) provide semantic guidance that helps EAs escape local optima, or (iii) encode problem structure that would otherwise be hard to capture with hand-designed heuristics. Liu et al. [24] showed that an LLM can be prompted to perform selection, crossover, and mutation directly, demonstrating that LLMs can function as zero-shot evolutionary optimizers. Wang et al. [30] further extended this idea to multi-agent settings where specialized LLM agents collectively drive an evolutionary search process.
Wu et al. [33] provide a taxonomy of these approaches, distinguishing between LLM-enhanced EAs (using LLMs to support or improve evolutionary search), EA-enhanced LLMs (using evolutionary strategies to tune or guide LLMs), and tightly integrated hybrid systems. EvoGPT fits into the category of LLM-enhanced EAs. Rather than replacing the evolutionary loop, we use LLMs to generate a semantically diverse initial population and to provide guided injections when the search plateaus. This combination retains the strengths of traditional SBST while exploiting the generative and semantic capabilities of modern LLMs.
@@ II-D LLMs and EAs for Automated Test Generation
The integration of LLMs and EAs specifically for automated test generation has recently gained momentum, motivated by the complementary strengths of semantic reasoning from LLMs and systematic exploration from EAs. Xiao et al. [34] investigate three places where LLM assistance can be used in an EA for test generation: (i) generating the initial population of test suites, (ii) injecting additional test cases during the evolutionary process to target uncovered code, and (iii) introducing new candidates when the search stagnates. Their experimental results show that LLM-based assistance yields the strongest benefits during the initialization phase and when applied after moderate stagnation, whereas overly frequent intervention provides diminishing returns.
Another influential system in this space is CodaMosa by Microsoft Research [23]. CodaMosa is an EA-based software test generation system that uses an LLM to escape structural coverage plateaus. Rather than relying solely on mutation and crossover, CodaMosa periodically queries an LLM to synthesize new tests aimed towards areas of the tested program that are less covered. We implemented a similar plateau-escaping mechanism in our EvoGPT.
Yang et al. [35] created pytLMtester, a test generation system that extends DynaMosa [28] for dynamically typed languages using LLMs. They run an EA to generate tests and use LLMs to infer object types, repair failed tests, and generate mutations. Hybrid methods for automated test generation have also been explored in other domains. SearchSYS [14] uses LLMs to generate high-level test templates that guide a search-based fuzzer in system-level simulators, showing that LLM-derived structure can improve bug-finding effectiveness.
@@ III The EvoGPT System
The initial population generation process instantiates five distinct LLM agent configurations, each defined by a different prompting strategy (e.g., focusing on edge cases, deep object chains, or creative assertion styles) as well as different temperature values. Each agent is queried asynchronously five times, resulting in a total of 25 diverse and structurally unique test suites. During this generation process, a lightweight repair loop inspired by TestART [18], is applied to each test suite: when a generated test fails due to compilation or runtime errors, the system captures the stack trace and attempts to fix it. This helps improve test validity and robustness early in the pipeline. Successful test suites go through an additional coverage enhancement step, which involves identifying uncovered branches in the code and calling an LLM to attempt to add tests that cover them. The resulting 25 test suites are passed to EvoGPT’s population optimization process, in which an EA iteratively evolves this population by applying selection, crossover, and mutation operations guided by a fitness function. Inspired by CodaMosa [23], when EvoGPT detects stagnation in the EA, it injects additional tests to help escape local optima. Here as well, multiple LLM agents are used to generate these injected tests, to encourage diversity. Next, we describe the details of the EvoGPT main processes: initial population generation process and population optimization.
@@ III-A Initial Population Generation
Prior to starting the initial population generation process, EvoGPT performs a lightweight pre-processing on the source code to reduce token complexity and context window. Specifically, we remove excessive or lengthy inline comments, documentation blocks, and unreachable code segments that could overload the model’s context window and negatively affect generation quality [16]. Then, we generate 25 test suites for a given class by instantiating five LLM agents and having each agent generate five test suites.
Importantly, each LLM agent is configured with a different temperature setting and a corresponding system prompt. The temperature value is used to control the stochasticity of the language model output. Prior research has shown that varying temperature values leads to a broader range of generated outputs, exposing different behavioral patterns and increasing semantic variance across samples [2]. To complement this, each agent is given a custom system prompt, which directs it to approach the test suite generation task from a slightly different perspective (e.g., focusing on execution paths, edge values, or assertion robustness). Table I lists for each of the five types of LLM agents we used, their temperature value, a qualitative description of their prompt (column “Prompt objective”), and intended purpose (column “Purpose”). The exact system prompts used for each LLM agent are included in the provided code. This diversity promotes both semantic variability and robustness in the generated tests. We empirically verified that this multi-prompt configuration produces semantically diverse initial test suites. A detailed analysis of the diversity metrics and statistical tests is presented in Section IV.
@@
After generating the initial set of tests, each LLM agent performs a generation-repair loop where it iteratively refines its test case in response to runtime feedback such as compilation errors or failing assertions. Such a generation-repair process was shown to enhance test validity and robustness in prior LLM-based systems such as ChatUniTest [12], TestART [18] and TestPilot [29]. In our system, this loop leverages a simplified version of TestART’s repair strategy: when a generated test fails with a runtime error or stack trace, EvoGPT programmatically extracts the stack trace and invokes a minimal repair heuristic to remove or correct the failing statement. Our minimal repair heuristic directly fixes common runtime and compilation errors. For instance, when an import error is raised, the repair loop programmatically traverses through the source code’s repository in order to fetch the right import paths. If our minimal repair heuristic fails to fix the generated test, we call the LLM with a repair prompt along with the stack trace that define the runtime or compilation error. The exact repair prompt is also included in our code. Using a programmatic, targeted, template-based repairs before calling an LLM-based repair has proved to be more stable and cost-effective for improving test pass rates than only using an LLM-based repair [18]
This repair loop is repeated at most four times, bounding the accumulation of error traces and code context passed to the LLM, thereby limiting prompt growth and reducing the risk of context window saturation [16]. If errors in the tests persist, the test methods responsible for these errors are automatically removed from the test suite.
@@
Once a valid test suite is synthesized, it is compiled and executed using JaCoCo,222www.jacoco.org/jacoco a Java Code Coverage tool, which provides detailed line and branch coverage information. The resulting coverage report is parsed and passed to the next module: the Coverage Analysis Agent. This agent uses the same temperature as the original generation agent to maintain behavioral consistency. It is then instructed to generate additional complementary tests, aimed specifically at covering the missed branches and lines reported in the coverage report. This strategy was found valuable for enhancing code coverage, as demonstrated in TestART [18]. Finally, the tests produced by both the initial generation agent and its corresponding coverage enhancement process are merged into a single composite test suite. This forms one member of the initial population in our evolutionary algorithm.